# Search for transient gamma-ray emission from magnetar flares using Fermi-LAT

### v1

@@ Abstract
We search for transient gamma-ray emission in the energy range from 0.1-300 GeV using data from the Fermi-LAT telescope in coincidence with magnetar flares. For our analysis we look for coincidence with 15 distinct flares from 11 magnetars using two distinct time windows of 1 day and 15 days. For 14 of these flares from 10 magnetars, we do not see any statistically significant gamma-ray emission. However, we see two gamma-ray flares from one magnetar, namely 1E 1048.1-5937, with combined significance of , observed after about 10 days from the peak of the X-ray flare. However, this magnetar is located close to the galactic plane (with galactic latitude of -0.52°) and this signal could be caused by contamination due to diffuse flux from gamma-ray sources in the galactic plane.
@@ I Introduction
Magnetars are a distinct class of neutron stars distinguished by their extraordinarily strong magnetic fields, which can reach extremely high values up to G and X-ray luminosities between ergs/sec Woods and Thompson (2006); Mereghetti (2008); Kaspi and Beloborodov (2017); Esposito et al. (2021). The magnetar emission stems from the decay and instability associated with their ultra-strong magnetic fields Duncan and Thompson (1992); Paczynski (1992); Thompson and Duncan (1995). Observationally, they are associated with soft-gamma ray repeaters (SGRs) and anomalous X-ray pulsars (AXP) Mereghetti (2008). Recently, a very intense radio burst characteristic of fast radio bursts, has also been detected from a galactic magnetar (SGR 1935+2154), which is a soft gamma-ray repeater CHIME/FRB Collaboration et al. (2020), hinting towards an association between magnetars and FRBs. Magnetars are known to emit a diversity of X-ray transients including short bursts, large outbursts with energies between ergs, giant flares, bursts with quasi-periodic oscillations, etc. Olausen and Kaspi (2014). These outbursts are emitted on a wide range of time scales from a few seconds to years Olausen and Kaspi (2014); Esposito et al. (2021). In this work, we focus on searching for gamma-ray emission from magnetars in coincidence with these X-ray outbursts.
@@
Some theoretical models of gamma-ray emission from magnetars have previously been discussed in Refs. Cheng and Zhang (2001); Zhang and Cheng (2002); Takata et al. (2013). The first ever theoretical calculation of the high energy gamma-ray flux for seven AXPs was done in Cheng and Zhang (2001). This work showed that the gamma-ray radiation arises due to curvature radiation produced by acceleration of electron-positron pairs, which are produced from collisions of high energy photons in the outer gap and soft X-rays from the stellar surface. This work argued that gamma-rays produced by magnetars could be detected by Fermi-LAT. Subsequently, the gamma-ray flux during the quiescent magnetar phase was calculated for two SGRs, viz. SGR 1806-20 and SGR 1900+14 Zhang and Cheng (2002). This work showed that the integral flux of SGR 1900+14 could be detected by Fermi-LAT Zhang and Cheng (2002). A model of pulsed GeV gamma-ray emission from magnetars at the time of X-ray outbursts was also proposed in Takata et al. (2013). In this model, the X-ray outbursts were proposed to be triggered by crust fractures, which could excite the Alfven waves that propagate into the outer magnetosphere. The GeV gamma rays are then produced by curvature radiation at the outer magnetosphere Takata et al. (2013). However, no flux predictions for individual magnetars were made in this work. As discussed in Abdo et al. (2010), even null results stemming from gamma-ray emission from magnetars provide interesting constraints on curvature/synchrotron emission mechanism in magnetars at high latitudes.
@@
In this work, we search for gamma-ray emission using the Fermi Large Area Telescope (LAT) gamma-ray telescope. The LAT is one of the two instruments onboard this detector and is sensitive to high-energy gamma rays from a plethora of astrophysical sources (Ajello et al., 2020). This detector is a pair-conversion telescope with a field of view of about 2.4 sr that is sensitive to photons between the energy range of 30 MeV to around 300 GeV and has been operating continuously since 2008 (Atwood et al., 2009).
@@
Previous searches for gamma-ray emission from magnetars with Fermi-LAT have been carried out in Abdo et al. (2010); Li et al. (2017). The first work which looked for gamma-ray emission from 0.1-10 GeV, was done by the Fermi-LAT collaboration using 17 months of Fermi-LAT data from 13 magnetars Abdo et al. (2010). No statistically significant emission was detected from any of the magnetars. Stringent upper limits on the persistent emission between and erg were obtained Abdo et al. (2010). Then, a follow-up search using about six years of Fermi-LAT data was carried out in Li et al. (2017), which looked for gamma-ray emission in the range from 0.1-10 GeV from 20 magnetars. No significant emission from these magnetars was detected and stringent upper limits between and erg were obtained Li et al. (2017). In addition to galactic magnetars, the Fermi-LAT collaboration also found evidence for GeV emission from an extragalactic magnetar in the Sculptor galaxy Fermi-LAT Collaboration et al. (2021). Therefore with more than 12 years of data, a search for gamma-ray emission at the time of magnetar bursts using the latest data is timely.
@@
For this work, we follow the same methodology as our previous works on analysis using Fermi-LAT, which looked for gamma-ray emission from galaxy clusters and OJ 287 Manna and Desai (2024a); Pasumarti and Desai (2024). This paper is structured as follows. We introduce the use of easyFermi, a software tool that largely simplifies the analysis of Fermi LAT data in Sect. II. We, then delve into the individual analysis of magnetar outbursts using dynamic time ranges and observe the obtained spectral energy distributions and light curves in Sect. III. We conclude in Sect. IV.
@@ II Observations and Data Analysis
For this work, we have obtained the time ranges of magnetar bursts observed in X-rays from the Magnetar Outburst Online Catalog111Magnetar Outburst Online Catalog : http://magnetars.ice.csic.es/#/outbursts Coti Zelati et al. (2018). The peak X-ray luminosity during these outbursts is between ergs. We have looked for a gamma-ray signal using 15 individual outbursts from 11 different magnetars. Among these 11 magnetars, 7 are SGRs and 4 AXPs. The McGill magnetar catalog Olausen and Kaspi (2014) on the other hand lists 30 known magnetars. However, for the remaining 19 magnetars, the outbursts (if any) were detected before the launch of the Fermi-LAT satellite in 2008, and hence were not considered in this work.
@@
The analysis of the Fermi-LAT data is done using easyFermi de Menezes (2022), which largely simplifies the data collection for analysis. The data files involving the spacecraft and photon files can either be downloaded directly from Fermi-LAT Data Query or can be automatically fetched by easyFermi, by specifying the magnetar’s J2000 coordinates and the time duration. The instrument response function used in the analysis is the Pass 8 Source v3. easyFermi also provides an option for downloading the required galactic (gll_iem_v07.fits) and isotropic (iso_P8R3_SOURCE_V3_v1.txt) diffuse emission models. For this work, we consider gamma-ray photons within an energy range of 0.1-300 GeV and a Region of Interest (RoI) of 15°. To minimize gamma-ray contamination from Earth’s limb, we adopt the standard zenith angle cut of 90°. The maximum likelihood analysis was performed on a square region with a spatial bin size of and 8 energy bins per decade. The spectral-spatial model includes all the cataloged sources lying in a region larger than the extent of the RoI. The magnetars are included in this model as point sources with a power-law spectrum, using their positions as listed in the McGill Online Magnetar Catalog. Source fitting is done using fit from fermipy with NewMinuit as the optimizer. All spectral parameters are left free for sources within of the RoI center, while the spectral parameters of the other sources are left to their catalog values. In all analyses, we have ensured a fit quality of 3, which is indicative of an accurate fit. The analysis uses the data from the 3rd data release of LAT data2224FGL-DR3: https://fermi.gsfc.nasa.gov/ssc/data/access/lat/12yr_catalog/, which covers an extended 12 year time range Abdollahi et al. (2020, 2022).
@@
The analysis, carried out by easyFermi, involves a binned likelihood analysis using the GTAnalysis class in fermipy Wood et al. (2017). For this purpose, a maximum likelihood test statistic, TS is used to evaluate the credibility of the gamma-ray source i.e. to quantify the likelihood of a source being real. The test statistic is defined as twice the maximum log-likelihood difference between an alternate and null hypothesis of no signal or rather, Mattox et al. (1996), where would signify the maximum likelihood value for a model with an additional test source (the ‘alternate’ hypothesis) and signifies the maximum likelihood value for a model without an additional source (the ‘null’ hypothesis). The same statistic is also used in neutrino astronomy Pasumarti and Desai (2022) and soft gamma-ray astronomy Manna and Desai (2024b). A higher TS value for a spatial bin indicates that the alternate hypothesis is valid, or in simpler terms, it is more likely that the observed gamma-ray flux in that spatial bin is due to a legitimate source. The significance of a detection is given by according to Wilk’s theorem Wilks (1938). We use a TS threshold value of 25 to quantify a detection.
| Source | Observed Peak | Peak Date | TS | TS | Upper Limits | 4FGL srcs |
| (MJD) | (Calendar) | (1 month) | (2 day) | within 3° | ||
| Swift J1822.3-1606 | 55758 | 2011-Jul-16 | 0.00 | 6.50 | 27 | |
| Swift J1834.9-0846 | 55780 | 2011-Aug-07 | 0.00 | 0.00 | 21 | |
| 1E 1048.1-5937* | 55926 | 2011-Dec-31 | 25.56 | 0.00 | 22 | |
| 1E 1048.1-5937* | 57592 | 2016-Jul-23 | 0.93 | 0.00 | 22 | |
| 1E 1048.1-5937* | 58120 | 2018-Jan-02 | 3.96 | 0.00 | 22 | |
| 1E 1547.0-5408* | 54742 | 2008-Oct-03 | 0.00 | 0.00 | 14 | |
| 1E 1547.0-5408* | 54854 | 2009-Jan-23 | 0.00 | 0.00 | 14 | |
| 1E 2259+586* | 56045 | 2012-Apr-28 | 0.00 | 0.00 | 13 | |
| CXOU J1647-4552* | 55829 | 2011-Sep-25 | 7.14 | 0.82 | 32 | |
| SGR 1935+2154 | 57075 | 2015-Feb-22 | 0.00 | 0.00 | 8 | |
| SGR 1935+2154 | 57526 | 2016-May-18 | 0.00 | 0.00 | 8 | |
| SGR 0418+5729 | 54993 | 2009-Jun-11 | 0.00 | 0.00 | 2 | |
| SGR 1833-0832 | 55275 | 2010-Mar-20 | 0.00 | 0.00 | 22 | |
| SGR 0501+4516 | 54701 | 2008-Aug-23 | 3.65 | 0.00 | 5 | |
| SGR 1745-2900 | 56411 | 2013-Apr-29 | 0.00 | 0.00 | 50 |
@@
The easyFermi analyses results also include spectral energy distributions (SEDs) and light curves. The former is computed using the sed method from fermipy, which fits the flux normalizations of the target source in each individual energy bin using a power law model with a fixed spectral index of 2.0. In our analyses, we have used 12 energy bins to plot the SEDs. In specific cases, though, we use fewer energy bins to better observe a weaker signal. The light curves of the magnetar point source are generated using the lightcurve method from fermipy, which fits the source characteristics through a series of time bins. We have adopted a standard value of 15 time bins for the light curves. For each magnetar in Table 1, we run the analysis in two distinct time windows (1 day and 15 days), centered around the peak outburst date. In the following sections, we shall go through the transient analysis results for each of the magnetars in detail.
@@ III Results
We now present our results for gamma-ray emission during the magnetar outbursts. For almost all the magnetars, we find that there is no conclusive evidence for gamma-ray emission coinciding with the onset of the magnetar outburst. For these magnetars, the TS values were much lesser than 25. Consequently, in such cases, we report the energy flux upper limits obtained through the likelihood analysis. Any signals in the vicinity of the magnetar were due to sources in the 4FGL-DR3 point source catalog Abdollahi et al. (2020, 2022). The only exception was for the magnetar 1E 1048-5937, when we observed a TS value for the first outburst at MJD of 55926, which corresponds to at least a 5 detection based on a 1 month time window. We now discuss the results for 1E 1048-5937 followed by the other magnetars. A tabular summary can be found in Table 1.
@@
1E 1048.1-5937 is an AXP, which is the closest magnetar to the Earth, at around 2900 parsecs away. A long-lasting X-ray flare was detected by SWIFT-XRT between 0.5-10 keV, beginning at 55926 MJD and having a peak flux of and a decay time of 550 days Archibald et al. (2015, 2020). The Fermi-LAT analysis of 1E 1048.1-5937’s outburst at 55296 MJD yielded a high TS value of 25.56 over a 1 month time window, corresponding to a 5 detection. However, the TS drops to zero during the 2 day window centered on the magnetar X-ray flare. This suggests that the gamma-ray emission during the larger, 1 month period is not exactly coincident with the peak in the X-ray emission.
@@
The results of the analyses can be summarized efficiently by a spatial significance map. The maps cover a square region with a spatial bin size of . These significance maps are present in Fig.1 and Fig.7. In Fig.1, we show two significance maps. The left panel presents the significance map derived by excluding 1E 1048.1-5937 from the model, effectively highlighting the -ray emission attributable to the magnetar. In contrast, the right panel depicts the residual significance map obtained after incorporating the magnetar into the region of interest (RoI) model, thereby representing the residual -ray emission after accounting for the contribution from 1E 1048.1-5937. There is a very visible peak in the significance maps, around a 5 detection, coinciding with the magnetar’s first X-ray outburst at 55926 MJD. Additionally, we modeled the magnetar as an extended source using a radial disk model. The results indicate no significant statistical evidence supporting the emission being attributed to an extended source, as the extension hypothesis test statistic, is negligible. The extension fit results are provided in Fig.2. The spectral energy distributions (SED) for the first outburst are presented in Fig. 3. The SED, obtained from the analysis of the one month time period, shows detected flux points in two energy bins at 500 and 1850 MeV.
We also look at the -ray light curves of 1E 1048.1-5937 to get a better idea on the duration of the gamma-ray signal. The -ray light curves are provided in Fig. 4 and Fig. 5 over a one-month and 5-day period, respectively. The light curves show detections, though they do not precisely align with the timing of the peak of the X-ray outburst at 55926 MJD. A rough estimate of the maximum duration of the -ray flare can be thought of as the total sum of the bin widths in which the detections are observed, from the month-long light curve. The estimated duration of the -ray flare, isolated to the two detected bins, is calculated to be 4.13 days, from Fig. 4. To achieve higher temporal resolution in the -ray light curve, we perform the LAT analysis over the shorter interval, focusing on the two detected bins. However, increasing the number of bins decreases the statistical significance. Consequently, we opt for a light curve spanning a 5-day period, divided into 17 bins to balance resolution with detection reliability. This -ray energy light curve of 1E 1048.1-5937 (refer Fig.5), over the sub-time interval of 5 days (07-Jan-2012 to 12-Jan-2012), shows two time bins which are not upper limits. The ranges of the two time bins are 55935.64 - 55935.93 MJD and 55937.10 - 55937.40 MJD respectively lasting about seven hours each, with TS values of 9.56 and 14.05 respectively. This corresponds to a 3.1 and 3.8 detection in the respective time bins. The detected time bins in the -ray light curve show an energy flux of and respectively. The two time bins are separated by about 1.17 days, which is a considerable gap. Therefore the signal consists of two distinct flares of lasting about seven hours instead of a single, contiguous flare. Further, we overlay the X-ray and Fermi-LAT -ray light curves for 1E 1048.1-5937. In order to directly compare the gamma-ray energy light curves from our Fermi LAT analysis to the X-Ray luminosity curves available on MOOC333Magnetar Outburst Online Catalog : http://magnetars.ice.csic.es/#/outbursts, we plot both in the same figure, to check if there is a direct correlation between the two. This plot showing both the fluxes can be found in Fig. 6.
@@
In addition to the observed burst at MJD 55926, two other X-ray flares were detected for this magnetar at MJD of 57592 and 58120. However, no gamma-ray emission was detected corresponding to these two flares. The spatial significance maps for the latter two, smaller X-ray outbursts at MJD of 57592 and 58120 are provided in Fig. 7. We can clearly see that the TS value is much less than 9 () and no gamma-ray signal is detected in coincidence with these flares.
@@
Therefore, for the first X-ray outburst at 55926, we find a 5 gamma-ray signal from 1E 1048.1-5937 about 10 days after the peak of the X-ray flare. The signal consists of two distinct gamma-ray flares, separated by around 1.17 days, with each flare lasting about seven hours. No gamma-ray signal is seen corresponding to the other two X-ray outbursts for this magnetar. We also note that a large number Fermi-LAT point sources from 4FGL-DR3 catalog are close to the magnetar. These include 4FGL J1046.7-6010, 4FGL J1048.5-5923, 4FGL J1054.0-5938, 4FGL J1054.7-6008c, 4FGL J1045.1-5940 and 4FGL J1055.7-6028, which are all within a 1°offset. A detailed characterization of the transient emission from these other point sources must also be done before concluding that the gamma-ray signal is definitively due to 1E 1048.1-5937. Furthermore, we caution that this magnetar is located close to the galactic plane where diffuse emission is strong and could contaminate observed signal Abdo et al. (2010).

### v2

@@ Abstract
We search for transient gamma-ray emission in the energy range from 0.1-10 GeV using data from the Fermi-LAT telescope in coincidence with magnetar flares. For our analysis, we look for coincidence with 15 distinct flares from 11 magnetars using three time windows of 1 day, 7 days, and 15 days. For 14 of these flares from 10 magnetars, we do not see any statistically significant gamma-ray emission. However, we see one possible gamma-ray flare from one magnetar, namely 1E 1048.1-5937, with combined significance of , (after considering the look-elsewhere effect) observed after about 10 days from the start of the X-ray flare. However, this magnetar is located close to the galactic plane (with galactic latitude of -0.52°), and this signal could be caused by contamination due to diffuse flux from gamma-ray sources in the galactic plane.
@@ I Introduction
Magnetars are a distinct class of neutron stars distinguished by their extraordinarily strong magnetic fields, which can reach extremely high values up to G and X-ray luminosities between ergs/sec Woods and Thompson (2006); Mereghetti (2008); Kaspi and Beloborodov (2017); Esposito et al. (2021); Rea and De Grandis (2025). The magnetar emission stems from the decay and instability associated with their ultra-strong magnetic fields Duncan and Thompson (1992); Paczynski (1992); Thompson and Duncan (1995). Observationally, they are associated with soft-gamma ray repeaters (SGRs) and anomalous X-ray pulsars (AXP) Mereghetti (2008). Recently, a very intense radio burst characteristic of fast radio bursts, has also been detected from a galactic magnetar (SGR 1935+2154), which is a soft gamma-ray repeater CHIME/FRB Collaboration et al. (2020), hinting towards an association between magnetars and FRBs. Magnetars are known to emit a diversity of X-ray transients including short bursts, large outbursts with energies between ergs, giant flares, bursts with quasi-periodic oscillations, etc. Olausen and Kaspi (2014). These outbursts are emitted on a wide range of time scales from a few seconds to years Olausen and Kaspi (2014); Esposito et al. (2021). In this work, we focus on searching for gamma-ray emission from magnetars in coincidence with these X-ray outbursts.
@@
Some theoretical models of gamma-ray emission from magnetars have previously been discussed in Refs. Cheng and Zhang (2001); Zhang and Cheng (2002); Takata et al. (2013). The first ever theoretical calculation of the high energy gamma-ray flux for seven AXPs was done in Cheng and Zhang (2001). This work showed that the gamma-ray radiation arises due to curvature radiation produced by acceleration of electron-positron pairs, which are produced from collisions of high energy photons in the outer gap and soft X-rays from the stellar surface. This work argued that gamma-rays produced by magnetars could be detected by Fermi-LAT. Subsequently, the gamma-ray flux during the quiescent magnetar phase was calculated for two SGRs, viz. SGR 1806-20 and SGR 1900+14 Zhang and Cheng (2002), and shown to be detectable for SGR 1900+14 by Fermi-LAT Zhang and Cheng (2002). A model of pulsed GeV gamma-ray emission from magnetars at the time of X-ray outbursts was also proposed in Takata et al. (2013). In this model, the X-ray outbursts were proposed to be triggered by crust fractures, which could excite the Alfven waves that propagate into the outer magnetosphere. The GeV gamma rays are then produced by curvature radiation at the outer magnetosphere Takata et al. (2013). However, no flux predictions for individual magnetars were made in this work. As discussed in Abdo et al. (2010a), even null results stemming from gamma-ray emission from magnetars provide interesting constraints on curvature/synchrotron emission mechanism in magnetars at high latitudes.
@@
In this work, we search for gamma-ray emission using the Fermi Large Area Telescope (LAT). The LAT is one of the two instruments onboard this detector and is sensitive to high-energy gamma rays from a plethora of astrophysical sources (Ajello et al., 2020). This detector is a pair-conversion telescope with a field of view of about 2.4 sr that is sensitive to photons between the energy range of 30 MeV to around 300 GeV and has been operating continuously since 2008 (Atwood et al., 2009).
@@
Previous searches for gamma-ray emission from magnetars with Fermi-LAT have been carried out in Abdo et al. (2010a); Li et al. (2017). The first work which looked for gamma-ray emission from 0.1-10 GeV, was done by the Fermi-LAT collaboration using 17 months of Fermi-LAT data from 13 magnetars Abdo et al. (2010a). No statistically significant emission was detected from any of the magnetars. Stringent upper limits on the persistent emission between and erg were obtained Abdo et al. (2010a). Then, a follow-up search using about six years of Fermi-LAT data was carried out in Li et al. (2017), which looked for gamma-ray emission in the range from 0.1-10 GeV from 20 magnetars. No significant emission from these magnetars was detected and stringent upper limits between and erg were obtained Li et al. (2017). In addition to galactic magnetars, the Fermi-LAT collaboration also found evidence for GeV emission from an extragalactic magnetar in the Sculptor galaxy Fermi-LAT Collaboration et al. (2021). Therefore, with more than 12 years of data, a search for gamma-ray emission in coincidence with magnetar outbursts using the latest data is timely, since the previous search for gamma-ray emission was only done in 2017.
@@
For this work, we follow the same methodology as our previous works on analysis using Fermi-LAT, which looked for gamma-ray emission from galaxy clusters and OJ 287 Manna and Desai (2024a); Pasumarti and Desai (2024). This paper is structured as follows. We introduce the use of easyFermi, a software tool that largely simplifies the analysis of Fermi LAT data in Sect. II. We then into the individual analysis of magnetar outbursts using dynamic time ranges and observe the obtained spectral energy distributions and light curves in Sect. III. We conclude in Sect. IV.
@@ II Observations and Data Analysis
For this work, we have obtained the time ranges of magnetar bursts observed in X-rays from the Magnetar Outburst Online Catalog111Magnetar Outburst Online Catalog : http://magnetars.ice.csic.es/#/outbursts Coti Zelati et al. (2018). The peak X-ray luminosity during these outbursts is between ergs/sec. We have looked for a gamma-ray signal using 15 individual outbursts from 11 different magnetars. Among these 11 magnetars, 7 are SGRs and 4 AXPs. The McGill magnetar catalog Olausen and Kaspi (2014), on the other hand, lists 30 known magnetars. However, for the remaining 19 magnetars, the outbursts (if any) were detected before the launch of the Fermi-LAT satellite in 2008, and hence were not considered in this work.
The durations of the magnetar flares are listed in Table 1. We also note that the observed peak of the magnetar flare (as in MOOC) is almost always coincident with the onset of the burst, based on the first available datum in the X-ray light curve, with the exceptions of SGR 0418+5729, SGR 1833-0832, and 1E 1048.1-5937 (corresponding to the flare at 57592 MJD), where the observed peak of the flare occurred 8 days, 5 days, and 6 days, respectively, after the onset.
@@
The analysis of the Fermi-LAT data is done using easyFermi de Menezes (2022), which largely simplifies the data collection for analysis. The data files involving the spacecraft and photon files can either be downloaded directly from Fermi-LAT Data Query or can be automatically fetched using easyFermi, by specifying the magnetar’s J2000 coordinates and the time duration. This analysis uses data from the 3rd data release of LAT data2224FGL-DR3: https://fermi.gsfc.nasa.gov/ssc/data/access/lat/12yr_catalog/, which covers an extended 12 year time range Abdollahi et al. (2020, 2022). The Pass 8 Source v3 instrument response functions (IRFs) were used in the analysis. easyFermi also provides an option for downloading the required galactic (gll_iem_v07.fits) and isotropic (iso_P8R3_SOURCE_V3_v1.txt) diffuse emission models.
For this work, we consider gamma-ray photons within the standard energy range of 0.1-10 GeV and a circular region of interest (RoI) of 10.6°. To minimize the gamma-ray contamination from Earth’s limb, we adopt the standard zenith angle cut of 90°. The maximum likelihood analysis was performed on a square region with a spatial bin size of and eight energy bins per decade, similar to Abdo et al. (2010a); Li et al. (2017). The spectral-spatial model includes all the cataloged sources lying in the square region larger than the extent of the region. The magnetars are included in this model as point sources with a power-law spectrum, using their positions as listed in the McGill Online Magnetar Catalog. Source fitting is done using fit from fermipy with NewMinuit as the optimizer. All spectral parameters are left free for sources within of the RoI center, while the spectral parameters of the other sources are left to their catalog values. The magnetars in Table 1 are modeled with a power law spectrum with all spectral parameters allowed to vary. In all analyses, we ensure a fit quality of 3, which is indicative of an accurate fit.
@@
The analysis, carried out by easyFermi, involves a binned likelihood analysis using the GTAnalysis class in fermipy Wood et al. (2017). For this purpose, a maximum likelihood test statistic, TS is used to evaluate the credibility of the gamma-ray source i.e. to quantify the likelihood of a source being real. The test statistic is defined as twice the maximum log-likelihood difference between an alternate and null hypothesis of no signal, or rather, Mattox et al. (1996), where would signify the maximum likelihood value for a model with an additional test source (the ‘alternate’ hypothesis) and signifies the maximum likelihood value for a model without an additional source (the ‘null’ hypothesis). The same statistic is also used in neutrino astronomy Pasumarti and Desai (2022) and soft gamma-ray astronomy Manna and Desai (2024b). A higher TS value for a spatial bin indicates that the alternate hypothesis is valid, or in simpler terms, it is more likely that the observed gamma-ray flux in that spatial bin is due to a legitimate source. For a source consisting of four free parameters (two for position and two for flux and spectral index), the distribution of TS is close to 1/2 of distribution with four degrees of freedom Abdo et al. (2010b). Therefore in this case a TS of 25 would correspond to a -value of or 4.1. For our analysis, since the two coordinates of the source are known, we have two degrees of freedom. Therefore, a TS value of 25, then corresponds to a statistical significance of Abdo et al. (2010b).
| Source | Flare Start | Flare Start Date | Outburst | TS | TS | 0.1 - 10 GeV | 4FGL srcs |
| (MJD) | (Calendar) | Timescale | ( days) | ( day) | within 5° | ||
| Swift J1822.3-1606 | 55758 | 2011-Jul-16 | 219 days | 0.00 | 6.45 | 56 | |
| Swift J1834.9-0846 | 55780 | 2011-Aug-07 | 97 days | 0.00 | 0.00 | 51 | |
| 1E 1048.1-5937* | 55926 | 2011-Dec-31 | 949 days | 24.91 | 0.00 | 45 | |
| 1E 1048.1-5937* | 57592 | 2016-Jul-23 | 211 days | 2.51 | 0.00 | 45 | |
| 1E 1048.1-5937* | 58120 | 2018-Jan-02 | 140 days | 0.00 | 0.00 | 45 | |
| 1E 1547.0-5408* | 54742 | 2008-Oct-03 | 23 days | 0.00 | 0.00 | 29 | |
| 1E 1547.0-5408* | 54854 | 2009-Jan-23 | 640 days | 0.00 | 0.00 | 29 | |
| 1E 2259+586* | 56045 | 2012-Apr-28 | 1298 days | 4.21 | 0.00 | 27 | |
| CXOU J1647-4552* | 55829 | 2011-Sep-25 | 28 days | 1.27 | 2.37 | 77 | |
| SGR 1935+2154 | 57075 | 2015-Feb-22 | 302 days | 0.00 | 0.00 | 21 | |
| SGR 1935+2154 | 57526 | 2016-May-18 | 71 days | 0.00 | 0.00 | 21 | |
| SGR 0418+5729 | 54993 | 2009-Jun-11 | 637 days | 0.00 | 0.00 | 9 | |
| SGR 1833-0832 | 55275 | 2010-Mar-20 | 159 days | 0.00 | 0.00 | 52 | |
| SGR 0501+4516 | 54701 | 2008-Aug-23 | 373 days | 3.48 | 0.00 | 6 | |
| SGR 1745-2900 | 56411 | 2013-Apr-29 | 1177 days | 8.30 | 0.00 | 116 |
@@
The results of the easyFermi analyzes also include spectral energy distributions (SEDs) and light curves. The former is computed using the sed method from fermipy, which fits the flux normalizations of the target source in each individual energy bin. In our analyses, we have used 12 energy bins to plot the SEDs. In specific cases, though, we use fewer energy bins to better observe a weaker signal. The light curves of the magnetar point sources are generated using the lightcurve method from fermipy, which fits the source characteristics through a series of time bins. We have adopted a standard value of 15 time bins for the light curves. For each magnetar in Table 1, we present the results for the analysis in two distinct time windows (1 day and 15 days), centered around the onset of the magnetar outburst. Similar to Fermi-LAT Collaboration et al. (2021), we also did the analysis in 7 day window, but since no additional signal was seen, these results are not reported. In the following sections, we shall go through the transient analysis results for each of the magnetars in detail.
@@ III Results
We now present our results for gamma-ray emission during the magnetar outbursts. For almost all the magnetars, we find that there is no conclusive evidence for gamma-ray emission coinciding with the onset of the magnetar outburst. For these magnetars, the TS values were much less than 25. Any signals in the vicinity of the magnetar were due to sources in the 4FGL-DR3 point source catalog Abdollahi et al. (2020, 2022). Consequently, in such cases, we report the energy flux upper limits obtained through the likelihood analysis. The 95% confidence level energy flux upper limits (0.1 - 10 GeV energy range) for magnetars, are then calculated using a power-law spectrum, with a fixed photon index of 2.5, as in Abdo et al. (2010a) and Li et al. (2017).
The only exception was for the magnetar 1E 1048-5937, wherein we observe a TS value for the first outburst at MJD of 55926, which has a significance of , based on a one month time window and incorporating the look-elsewhere effect, considering the fact that we have considered three distinct temporal windows. We now discuss the results for 1E 1048-5937 followed by the other magnetars. A tabular summary of all our results can be found in Table 1.
@@
1E 1048.1-5937 is an Anomalous X-ray Pulsar (AXP), at around 9 kiloparsecs away. A long-lasting X-ray flare was detected by SWIFT-XRT between 0.5-10 keV, beginning at 55926 MJD and having a peak flux of and a decay time of 550 days Archibald et al. (2015, 2020). The Fermi-LAT analysis of 1E 1048.1-5937’s outburst at 55926 MJD yielded a high TS value of 24.91 over a 1-month time window. However, the TS value drops to zero during the 2 day window centered on the magnetar X-ray flare. This suggests that the gamma-ray emission during the larger, one month period is not exactly coincident with the peak in the X-ray emission.
@@
The results of the analyses can be summarized efficiently using a spatial significance map. The maps cover a square region with a spatial bin size of . These significance maps can be found in Fig. 1. In Fig. 1, we show two significance maps. The left panel presents the significance map derived by excluding 1E 1048.1-5937 from the model, effectively highlighting the -ray emission attributable to the magnetar. In contrast, the right panel depicts the residual significance map obtained after incorporating the magnetar into the region of interest (RoI) model, thereby representing the residual -ray emission after accounting for the contribution from 1E 1048.1-5937. There is a very visible peak in the significance maps, (with significance of ), coinciding with the magnetar’s first X-ray outburst at 55926 MJD.
Upon performing a spatial localization for the magnetar point source, the best-fit position is given by (ra:162.424∘, dec:-60.000∘), with a 99% positional uncertainty radius of 0.22∘. We also note that the maxima in TS lies within this region, as illustrated in Fig. 1.
Additionally, we tested the incorporation of a new unidentified gamma-ray source exactly coinciding positionally with the maxima in TS, along with the magnetar point source. On performing a spatial localization for the unidentified source, we determine that the best-fit position is (ra:162.412∘, dec:-60.013∘), with a 99% positional uncertainty radius of 0.22∘. Notably, 1E 1048.1-5937 lies well within this region and furthermore, no existing Fermi-LAT sources are found within this region in the 4FGL-DR3 or 4FGL-DR4 LAT catalogs Abdollahi et al. (2020); Ballet et al. (2023) within the 99% positional uncertainty.
We also looked for other possible multi-wavelength candidate counterparts, using SIMBAD, within the 99% positional uncertainty radius of 0.22∘, which corresponds to a 13.2 arc-minute radius. We find a total of 345 sources, including one high-mass X-ray binary (2E-2336) at an angular separation of 13.58 arc-seconds and five X-ray point sources with angular separations ranging from arc-seconds, from the magnetar. Other sources include main-sequence stars, eclipsing binary stars, long period variable stars and RR Lyrae variable stars. These sources (e.g. the high-mass X-ray binary 2E 2336) might be viable alternatives to the source of the observed gamma-ray emission.
Further, we modeled the magnetar as an extended source using a radial disk model. The results indicate that there is no significant statistical evidence supporting the emission being attributed to an extended source, as the extension hypothesis test statistic, is negligible.
The spectral energy distributions (SED) for the first outburst are presented in Fig. 2. The SED, obtained from the analysis of the one month time period, shows a detected flux point in the energy bin at 500 MeV.
We also look at the standard -ray light curve to get a better idea of the duration of the gamma-ray signal. We adopt a standard binning of 15 bins, over the 1-month time range. The -ray light curve is provided in Fig. 3 for the 1-month time range. In these light curves, all data points with are shown as upper limits, whereas if , the corresponding flux along with its Poisson error is shown.
Further, we overlay the X-ray and Fermi-LAT -ray light curves for 1E 1048.1-5937. In order to directly compare the gamma-ray energy light curves from our Fermi LAT analysis to the X-Ray luminosity curves available on MOOC333Magnetar Outburst Online Catalog : http://magnetars.ice.csic.es/#/outbursts, we plot both the X-ray and gamma-ray light curves in the same figure, to check if there is a direct correlation between the two. This plot showing both the fluxes can be found in Fig. 4. Therefore, with only one data point, it is hard to make a definitive statement about the correlation between the X-ray and gamma-ray light curve. We also note that the TS value of the first data point (with a bin-size of 23.2 days) in Fig. 4 is about 15.65 and is suppressed, compared to that obtained within a one month time window.
@@
In addition to the observed burst at MJD 55926, two other X-ray flares were detected for this magnetar at MJD of 57592 and 58120. The timescales of these two flares are also provided in Table 1. The duration of the flare corresponding to 1E 1048.1-5937’s third outburst in early 2018 is obtained from Archibald et al. (2020). However, no gamma-ray emission was detected corresponding to these two flares. The spatial significance maps for the latter two, smaller X-ray outbursts at MJD of 57592 and 58120 are provided in Fig. 5. We can clearly see that the TS value is much less than 9 and no gamma-ray signal is detected in coincidence with these flares.
@@
Therefore, for the first X-ray outburst at 55926, we find a gamma-ray signal from 1E 1048.1-5937 about 10 days after the peak of the X-ray flare. No gamma-ray signal is seen corresponding to the other two X-ray outbursts for this magnetar. Finally, we caution that this magnetar is located close to the galactic plane where diffuse emission is strong and could contaminate the observed signal Abdo et al. (2010a).